The Niagara River delivers an average of 5470 m3 of water per second to the top of Niagara Falls, where it drops 48.5 m. If all the potential energy of that water could be converted to electrical energy, how much electrical power could Niagara Falls generate? (Note: The mass of 1 m3 of water is 1000 kg).
1
Expert's answer
2019-07-23T15:54:59-0400

"P=\gamma\times Q\times H"

Where P is the power generated

"\gamma=" Specific weight= Density x aceleration due to gravity= 1000x9.81=9810 N/m3=9.81 KN/m3 which is approximately to 10KN/m3

Q= Discharge or volumetric flow rates and given as 5470m3/s

H = head or height of flow = 48.5m

Thus

P=10×5470×48.5

P=2,652,950 KN-m/s or KW